OpenStudy (anonymous):

Solve, using the substitution method. y = 4x + 6 5x – y = 6 A. There are an infinite number of solutions. B. There is no solution. C. (18, 12) D. (12, 54)

OpenStudy (anonymous):

y = 4x + 6 -6 -6 y - 6 = 4x y/4 - 6/4 = 4x/4 y/4 - 1.5 = x 5x – y = 6 5(y/4 - 1.5) - y = 6 Then you can just plug into https://www.desmos.com/calculator and y = 54 then substitute y into either equation and solve for x

OpenStudy (anonymous):

y = 4x + 6 54 = 4x + 6 -6 -6 48 = 4x 48/4 = 4x/4 12 = x So the solution is (12, 54)

OpenStudy (anonymous):

@kira12341234 are you there

OpenStudy (anonymous):

oh im so sorry i had to write a paper thank you so much for the help